OpenStudy (anonymous):

Why is 4/16 equivalent to 0.25?

OpenStudy (anonymous):

\[\frac{4}{16}=\frac{1}{4}=\frac{25}{100}=.25\]

OpenStudy (anonymous):

^ boom